# SALT: Parameter-Efficient Fine-Tuning via Singular Value Adaptation with Low-Rank Transformation

### v1

@@ 4 Results and Discussion
Ablation Study. We conducted comprehensive ablation experiments on the Drive dataset [22] to evaluate the impact of different architectural components and rank configurations in SALT. Fig 4(a) presents the results across various combinations of LoRA ranks (4, 64, 256) and SALT ranks for Multi-Head Attention (MHA), neck convolutional layers (NECK-Conv2d), and Multi-Layer Perceptron (MLP) modules. The SALT rank for a given layer refers to the number of dominant singular values in the layer’s weight matrix that are adaptively scaled and shifted during fine-tuning. Key findings are: 1) increasing the LoRA rank from 4 to 256 consistently improved Dice scores ($0.72\ \rightarrow 0.76$) while maintaining a modest parameter increase from 0.46% to 5.86% and 2) reducing MHA and MLP ranks from 700 to 200 further enhanced performance, increasing Dice scores from 0.75 to 0.76 with LoRA rank 256. These results suggest that higher LoRA ranks improve expressiveness, while component-specific rank selection is crucial for optimal performance. The optimal configuration (LoRA rank 256 and components rank 200) achieved a Dice score of 0.76 with only 5.84% of trainable parameters, demonstrating SALT’s effective balance of performance and efficiency.
@@ 5 Conclusion
In this paper, we introduce SALT, a novel PEFT framework that adapts the Segment Anything Model (SAM) to domain-specific tasks, particularly medical image segmentation. By utilizing SVD-based scaling and shifting of the top singular values with LoRA-based low-rank updates, SALT achieves efficient and flexible adaptation, enabling precise segmentation with minimal computational overhead. Experiments show that SALT outperforms state-of-the-art SAM PEFT adaptations while preserving parameter efficiency and adaptability. Future work will extend SALT to other foundation models and explore dynamic rank selection to enable optimal parameter allocation.
6 Acknowledgments
The authors gratefully acknowledges Juan Lugo for his contributions through out the project.

### v2

@@ 4.4 Ablation Study
We conducted comprehensive ablation experiments on the DRIVE dataset [Staal et al.(2004)Staal, Abramoff, Niemeijer, Viergever, and van Ginneken] to evaluate the impact of different architectural components and rank configurations in SALT. Fig. 4 presents the results across various combinations of LoRA ranks (4, 64, 256) and SALT ranks for MHA, neck convolutional layers (NECK-Conv2d), and MLP modules. The SALT rank for a given layer refers to the number of dominant singular values in the layer’s weight matrix that are adaptively scaled and shifted during fine-tuning. For example, a SALT rank of 200 for the Multi-Layer Perceptron modules means that for each weight matrix, we update the top 200 singular values using scale and shift parameters, while the rest are updated using the low-rank updates. Key findings are: 1) increasing the LoRA rank from 4 to 256 consistently improved Dice scores (0.720.760.72\ $\rightarrow$ 0.76) while maintaining a modest parameter increase from 0.46% to 5.86% and 2) reducing MHA and MLP ranks from 700 to 200 further enhanced performance, increasing Dice scores from 0.75 to 0.76 with LoRA rank 256. These results suggest that higher LoRA ranks improve expressiveness, while component-specific SALT rank selection is crucial for optimal performance. By decreasing the SALT rank, we allow less dominant singular values to be adapted through LoRA updates, which enhances the model’s adaptation capacity. The optimal configuration (LoRA rank 256 and components rank 200) achieved a Dice score of 0.76 with only 5.84% of trainable parameters, demonstrating SALT’s effective balance of performance and efficiency.
@@ 5 Conclusion
In this paper, we introduce SALT, a novel PEFT framework that adapts the Segment Anything Model (SAM) to domain-specific tasks, particularly medical image segmentation. By utilizing SVD-based scaling and shifting of the top singular values with LoRA-based low-rank updates, SALT achieves efficient and flexible adaptation, enabling precise segmentation with minimal computational overhead. Experiments show that SALT outperforms state-of-the-art SAM PEFT adaptations while preserving parameter efficiency and adaptability. While we have focused on SAM and its variants, many other large-scale vision backbones, Vision Transformers (ViTs), and even multi-modal encoders could benefit from SALT. In future work, applying SALT to these architectures will test its versatility across different embedding dimensions, attention patterns, and domain gaps. Moreover, the current design fixes the selection of ranks and SALT hyperparameters per layer. In future work, we aim to develop methods to automatically choose each layer’s rank to tailor parameter allocation to each layer’s intrinsic complexity. Finally, we will add boundary-aware training and small multi-scale tweaks to better capture fine edges and lower HD95, especially when data is scarce or ranks are very low.